(1) Simplify the square root: √49 *
Lesson 1
-7
7 <---- ?
49

(2) Classify the following: Pi (3.1415...) *
Lesson 1
Rational
Irrational <---?

(3) Solve for g: 3g + 4 = 19 *
Lesson 2
g = 5
g = 15
g = 45

(4) Solve for x: 9x - 2x = 28 *
Lesson 3
x = 28/11
x = 7
x = 4

(5) Solve for p: .4p - 6 = 32 *
Lesson 4
p = 2.4
p = 38
p = 95

(6) Write an equation to represent the following: Steve rented a car for $12.00 per day plus $.25 per mile driven. His total bill was $51.00. *
Lesson 5 - (2 points)

(7) Solve for y: 7y – 13 = 6y + 3 *
Lesson 6
y = 1
y = 16
y = 10

(8) Solve for k: -2k + 2 < 14 *
Lesson 7 - (2 points)
1
Expert's answer
2017-03-15T14:22:06-0400
(1) Simplify the square root: √49 *
Lesson 1
-7
7
49
Answer: ±7.

(2) Classify the following: Pi (3.1415...) *
Lesson 1
Rational
Irrational
Answer: Irrational.

(3) Solve for g: 3g + 4 = 19 *
Lesson 2
g = 5
g = 15
g = 45
Answer: g=5.

(4) Solve for x: 9x - 2x = 28 *
Lesson 3
x = 28/11
x = 7
x = 4
Answer: x=4.

(5) Solve for p: .4p - 6 = 32 *
Lesson 4
p = 2.4
p = 38
p = 95
Answer: p=9.5.

(6) Write an equation to represent the following: Steve rented a car for $12.00 per day plus $.25 per mile driven. His total bill was $51.00. *
Lesson 5 - (2 points)
x - mile.
Answer: 0.25x+12=51.

(7) Solve for y: 7y – 13 = 6y + 3 *
Lesson 6
y = 1
y = 16
y = 10
Answer: y=16.

(8) Solve for k: -2k + 2 < 14 *
Lesson 7 - (2 points)
Answer: k>-6.